# MarS-FM: Generative Modeling of Molecular Dynamics via Markov State Models

### v2

@@ C.3 JSD for pairwise RMSD and global RMSF, and Pearson $r$ for Folding Free Energies
Pairwise RMSD JSD Global RMSF JSD $\Delta G_{\text{fold}}~r$ MD (Oracle) $0.16$ $\pm~.005$ $0.09$ $\pm~.003$ $0.90$ $\pm~.002$ MDGen-100 $0.22$ $\pm~.008$ $0.31$ $\pm~.009$ $0.82$ $\pm~.006$ MDGen-20 $0.38$ $\pm~.005$ $0.52$ $\pm~.010$ $0.85$ $\pm~.005$ MDGen-100 (in parallel) $0.54$ $\pm~.004$ $0.58$ $\pm~.004$ $0.83$ $\pm~.004$ MDGen-20 (in parallel) $0.73$ $\pm~.002$ $0.71$ $\pm~.001$ $0.86$ $\pm~.001$ BioEmu $0.70$ $\pm~.001$ $0.72$ $\pm~.001$ $0.71$ $\pm~.002$ MarS-FM $\oplus$ MDGen-20 $\bf 0.15$ $\pm~.001$ $\bf 0.24$ $\pm~.001$ $\bf 0.88$ $\pm~.001$ MarS-FM $0.18$ $\pm~.001$ $0.28$ $\pm~.001$ $0.87$ $\pm~.003$
C.4 MD-Cath at 320K
@@
In contrast to the high-temperature replicas, trajectories at 320 K display no large-scale domain motions. Both the radius of gyration and secondary structure fractions exhibit nearly an order-of-magnitude lower variability (Table 6). Similarly, all-atom RMSF and secondary structure fluctuations remain tightly clustered around their native values. These statistics confirm that proteins largely remain stable at lower temperatures, with no significant conformational changes—consistent with the findings reported by Mirarchi et al. (2024).
@@
Appendix D Broader Impacts

### v3

@@ C.6 JSD for pairwise RMSD and global RMSF, and Pearson $r$ for Folding Free Energies
Pairwise RMSD JSD Global RMSF JSD $\Delta G_{\text{fold}}\penalty 10000\ r$ MD (Oracle) $0.16$ ± $\pm\penalty 10000\ .005$ $0.09$ ± $\pm\penalty 10000\ .003$ $0.90$ ± $\pm\penalty 10000\ .002$ MDGen-100 $0.22$ ± $\pm\penalty 10000\ .008$ $0.31$ ± $\pm\penalty 10000\ .009$ $0.82$ ± $\pm\penalty 10000\ .006$ MDGen-20 $0.38$ ± $\pm\penalty 10000\ .005$ $0.52$ ± $\pm\penalty 10000\ .010$ $0.85$ ± $\pm\penalty 10000\ .005$ MDGen-100 (in parallel) $0.54$ ± $\pm\penalty 10000\ .004$ $0.58$ ± $\pm\penalty 10000\ .004$ $0.83$ ± $\pm\penalty 10000\ .004$ MDGen-20 (in parallel) $0.73$ ± $\pm\penalty 10000\ .002$ $0.71$ ± $\pm\penalty 10000\ .001$ $0.86$ ± $\pm\penalty 10000\ .001$ BioEmu $0.70$ ± $\pm\penalty 10000\ .001$ $0.72$ ± $\pm\penalty 10000\ .001$ $0.71$ ± $\pm\penalty 10000\ .002$ MarS-FM $\oplus$ MDGen-20 $\bf 0.15$ ± $\pm\penalty 10000\ .001$ $\bf 0.24$ ± $\pm\penalty 10000\ .001$ $\bf 0.88$ ± $\pm\penalty 10000\ .001$ MarS-FM $0.18$ ± $\pm\penalty 10000\ .001$ $0.28$ ± $\pm\penalty 10000\ .001$ $0.87$ ± $\pm\penalty 10000\ .003$
C.7 Validity Metrics
We report validity metrics that directly probe native-contact preservation, Cα–Cα geometry, and backbone covalent structure on the MD-Cath 450 K test set. Following the EquiJump (costa2024equijump), we compute Jensen–Shannon divergences (JSDs) between MD and model-generated trajectories for fraction of native contacts (FNC) and for the Global Distance Test–Total Score (GDT-TS). We additionally report backbone bond- and angle-RMSZ values.
Fraction of native contacts.
FNC quantifies how well a sampled conformation preserves the interatomic contact pattern of a reference structure (best2013native). Given the set of native contacts $\mathcal{C}_{\mathrm{ref}}$ defined on the reference structure (pairs of atoms closer than a cutoff $d_{\text{cut}}$), the FNC of a conformation $x$ is
$\mathrm{FNC}(x)=\frac{\#\{(i,j)\in\mathcal{C}_{\mathrm{ref}}:d_{ij}(x)\leq d_{\text{cut}}\}}{\#\mathcal{C}_{\mathrm{ref}}}\,$, (11)
and we report the JSD between the FNC distributions of generated and MD trajectories.
Global Distance Test–Total Score.
GDT-TS measures Cα-level similarity to a reference structure (zemla2003lga). For a conformation $x$ with Cα coordinates $\{r_{i}(x)\}_{i=1}^{L}$ and reference coordinates $\{r_{i}^{\ast}\}_{i=1}^{L}$, we calculate,
$\mathrm{GDT\text{-}TS}(x)=\frac{1}{4}\sum_{d\in\{1,2,4,8\}\,\text{\AA }}\frac{\#\{i\in\{1,\dots,L\}:\lVert r_{i}(x)-r_{i}^{\ast}\rVert\leq d\}}{L}\,.$ (12)
We again compute the JSD between the GDT-TS distributions of generated and MD trajectories.
Backbone covalent-geometry RMSZ.
To assess bonding quality, we compute backbone bond-length and bond-angle root-mean-square $Z$-scores (RMSZ) (chen2010molprobity). For each backbone bond $k$, we form a $Z$-score
$z_{k}=\frac{\ell_{k}-\ell_{k}^{\mathrm{ideal}}}{\sigma_{k}}\,$, (13)
where $\ell_{k}$ is the observed bond length, and $\ell_{k}^{\mathrm{ideal}}$ and $\sigma_{k}$ are the Engh–Huber crystallographic target mean and standard deviation for that bond type (engh1991accurate). The bond RMSZ is then
$\mathrm{Bond\ RMSZ}=\sqrt{\frac{1}{N_{\mathrm{bond}}}\sum_{k=1}^{N_{\mathrm{bond}}}z_{k}^{2}}\,$, (14)
with an analogous definition for the angle RMSZ over backbone bond angles. Values close to the MD oracle indicate MD-like covalent geometry, whereas large RMSZ reflects systematic distortions.
Table 8 summarizes these additional validity metrics. Both MarS-FM and MarS-FM $\oplus$ MDGen-20 achieve FNC- and GDT-TS-based JSDs that are closest to MD, indicating that they best reproduce MD-like distributions of native contacts and Cα distances. For covalent geometry, the MDGen-20/100 parallel variants obtain the lowest RMSZ by construction, as they perform only minimal moves around a single conditioning structure; however, they underperform strongly on dynamical diversity metrics (Table 2). In contrast, MarS-FM models explore configuration space substantially more, yet still markedly improve bond and angle RMSZ over MDGen and BioEmu.
Table 8: JSD $\downarrow$ for fraction of native contacts (FNC) and GDT-TS, and RMSZ $\downarrow$ for backbone bond lengths and bond angles. Results are based on sampled trajectories of 500 conformations compared to ground-truth MD distributions and averaged over 5 inference runs.
FNC JSD GDT-TS JSD Bond RMSZ Angle RMSZ MD (Oracle) $0.13$ ± $\pm\penalty 10000\ .003$ $0.11$ ± $\pm\penalty 10000\ .004$ $2.23$ ± $\pm\penalty 10000\ .012$ $2.19$ ± $\pm\penalty 10000\ .009$ MDGen-100 $0.20$ ± $\pm\penalty 10000\ .004$ $0.17$ ± $\pm\penalty 10000\ .004$ $76.58$ ± $\pm\penalty 10000\ 2.041$ $5.27$ ± $\pm\penalty 10000\ .077$ MDGen-20 $0.30$ ± $\pm\penalty 10000\ .003$ $0.22$ ± $\pm\penalty 10000\ .002$ $407.82$ ± $\pm\penalty 10000\ 5.857$ $7.00$ ± $\pm\penalty 10000\ .031$ MDGen-100 (in parallel) $0.29$ ± $\pm\penalty 10000\ .001$ $0.31$ ± $\pm\penalty 10000\ .002$ $3.87$ ± $\pm\penalty 10000\ .009$ $2.56$ ± $\pm\penalty 10000\ .002$ MDGen-20 (in parallel) $0.40$ ± $\pm\penalty 10000\ .001$ $0.42$ ± $\pm\penalty 10000\ .001$ $2.88$ ± $\pm\penalty 10000\ .001$ $2.32$ ± $\pm\penalty 10000\ .001$ BioEmu $0.53$ ± $\pm\penalty 10000\ .001$ $0.49$ ± $\pm\penalty 10000\ .001$ $15.25$ ± $\pm\penalty 10000\ 3.151$ $2.90$ ± $\pm\penalty 10000\ .001$ MarS-FM $\oplus$ MDGen-20 $0.18$ ± $\pm\penalty 10000\ .001$ $0.16$ ± $\pm\penalty 10000\ .000$ $8.34$ ± $\pm\penalty 10000\ .121$ $2.68$ ± $\pm\penalty 10000\ .001$ MarS-FM $0.15$ ± $\pm\penalty 10000\ .001$ $0.12$ ± $\pm\penalty 10000\ .001$ $12.62$ ± $\pm\penalty 10000\ .144$ $3.06$ ± $\pm\penalty 10000\ .001$
C.8 MD-Cath at 320K
@@
In contrast to the high-temperature replicas, trajectories at 320 K display no large-scale domain motions. Both the radius of gyration and secondary structure fractions exhibit nearly an order-of-magnitude lower variability (Table 9). Similarly, all-atom RMSF and secondary structure fluctuations remain tightly clustered around their native values. These statistics confirm that proteins largely remain stable at lower temperatures, with no significant conformational changes—consistent with the findings reported by mirarchi2024mdcath.
@@
Appendix D LLM Usage
We used large language models (GPT-5) to assist in improving the grammar of the manuscript, as well as facilitate writing code.
Appendix E Broader Impacts